# Scale Reliant Inference

### v4

@@ 4.1 A Continuum between Scale Invariant and Scale Reliant Inference
Identifiability is not a dichotomous state but rather a continuum based on the identification region of a parameter. This implies that there is a spectrum between SRI and SII. Here, we define this progression before showing how these definitions apply to inferential problems described in existing literature.
@@
In the extreme case where ${\mbox{Id}}_{\theta}(P)$ is a singleton set, $\theta$ is identified, and the estimation problem is SII. At the opposite end, where ${\mbox{Id}}_{\theta}(P)=\Theta$, $\theta$ is fully unidentified, and the estimation problem is Full SRI. Full SRI represents the extreme case where, due to the arbitrary scale of the data, all values of $\theta\in\Theta$ are consistent with the observed measurements. In between these two cases, where ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$, $\theta$ is partially identified, and the estimation problem is Partial SRI.
@@
In Section 5, we show that the distinction between SII, Full SRI, and Partial SRI is directly related to the conditions under which certain types of inference (e.g., Frequentist versus Bayesian) are possible. Before providing those results, we first motivate that all three types of estimation problems are encountered within the literature. To do this, we describe the two relevant factors that determine whether a problem is SII, Full SRI, or Partial SRI: the degree to which the estimand depends on scale and the size of the identification region of $W^{{\tiny\mbox{tot}}}$ at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$.
4.1.1 A Characterization of Target Estimands
Consider the following three target estimands:
$\eta_{n}$ $=\log\frac{\mathbf{W}_{d_{i}n}}{\mathbf{W}_{d_{j}n}},$ (3)
$\theta_{d}$ $=\operatorname*{mean}_{n:x_{n}=1}\log\mathbf{W}_{dn}-\operatorname*{mean}_{n:x_{n}=0}\log\mathbf{W}_{dn},$ (4)
$\Sigma_{d_{i},d_{j}}$ $=\text{cov}(\log\mathbf{W}_{d_{i}\cdot},\log\mathbf{W}_{d_{j},\cdot}).$ (5)
The first is the type of scale invariant quantity purported by the field of CoDA. To see why it does not depend on scale consider that
@@ 4.1.1 A Characterization of Target Estimands
The second is the LFC estimand introduced in Section 2 while the third is a quantity of central interest to micro- and macro-ecological modeling (e.g., Friedman and Alm, (2012)). While both have been criticized in the CoDA literature as not being scale invariant (Lovell et al.,, 2015), these estimands differ in the degree to which they rely on scale. Next we show that LFC estimand is fully scale reliant while the covariance estimand is partially scale reliant.
We define the scale invariance/reliance of an estimand as follows. Let $\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$ denote the set of all systems that have a particular composition $\mathbf{W}^{{\tiny\mbox{comp}}}$. Let $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$) denote the image of this set with respect to the target estimand:
$\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})=\{\theta\in\Theta:\exists\,\,W^{{\tiny\mbox{tot}}}\in\mathcal{W}^{\tiny\mbox{tot}},\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})=\theta\},$
where $\mathcal{W}^{\tiny\mbox{tot}}$ is the space of scale vectors (e.g., $\mathbb{Z}^{N+}$ if $W$ is integer valued or $\mathbb{R}^{N+}$ if continuous). We define scale invariance/reliance based on the size of this image. If $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})$ is a singleton set, we say that $\theta$ is scale invariant at $\mathbf{W}^{{\tiny\mbox{comp}}}$; otherwise, it is scale reliant at $\mathbf{W}^{{\tiny\mbox{comp}}}$. We sub-categorize scale reliant estimands as fully scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))=\Theta$ or partially scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))\subsetneq\Theta$. We will drop the phrase at $\mathbf{W}^{{\tiny\mbox{comp}}}$ if an estimand is scale invariant or reliant for all $\mathbf{W}^{{\tiny\mbox{comp}}}\in\mathcal{W}^{\tiny\mbox{comp}}$ where $\mathcal{W}^{\tiny\mbox{comp}}$ is the space of compositional matrices. To summarize, using only compositional information $\mathbf{W}^{{\tiny\mbox{comp}}}$, we can learn everything about scale invariant estimands, something about partially scale reliant estimands, and nothing about fully scale reliant estimands.
It follows from Equation 6 that the log-ratio estimand is scale invariant based on this definition. Moreover, from Equation 2, it follows that the LFC estimand is fully scale reliant. In Supplementary Section B.1, we prove that the covariance estimand is partially scale reliant so long as the covariance is finite. Elsewhere, we have shown how common differential set analyses (e.g., gene set enrichment analysis (Subramanian et al.,, 2005)) correspond to estimands that are scale reliant for some $\mathbf{W}^{{\tiny\mbox{comp}}}\in\mathcal{W}^{\tiny\mbox{comp}}$ and scale invariant for others (McGovern et al.,, 2023). In short, there is a tremendous diversity of estimands that goes beyond whether they are scale invariant or not.
@@ 4.1.2 Characterization of Compositional Surveys
Subtlety is required when considering compositional surveys. When authors equate data with arbitrary scale with compositional data (Gloor et al.,, 2017)), they are assuming that the observed data provides no information about the system scale. This corresponds to an extreme from of compositional survey where ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)=\mathcal{W}^{\tiny\mbox{tot}}$. We call these full compositional surveys. In some cases (e.g., human gut microbiome surveys), this may be a reasonable model. Yet in other cases, we expect that the observed data provides important, yet partial, information about scale such that ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)\subsetneq\mathcal{W}^{\tiny\mbox{tot}}$. For example, consider the political polling scenario at the start of this article. If the $n$-th poll surveys 10,000 people, we can infer that $W^{{\tiny\mbox{tot}}}_{n}\geq 10,000$. We call these partial compositional surveys.
4.1.3 Relating Definitions back to SRI and Real Problems
In Supplementary Section B.2, we provide a number of theoretical results which establish how estimands and compositional surveys combine to determine if an estimation problem is Full SRI, Partial SRI, or SII. A particularly intuitive result is that, in the setting of a full compositional survey, if an estimand is fully scale reliant, partially scale reliant, or scale invariant, then the resultant estimation problem is Full SRI, Partial SRI, or SII, respectively.
To relate back to real problems addressed in existing literature, recall that calling the data compositional is effectively modeling the observed data as an observation of a full compositional survey. Within the analysis of human microbiota, we find examples of researchers calling these data compositional, yet they are trying to use these data to estimate all three of the estimands presented above (Fernandes et al.,, 2014; Friedman and Alm,, 2012; Silverman et al.,, 2017). In short, this distinction between SII, Partial SRI, and Full SRI are not simply theoretical constructs but relate to real problems encountered in the literature.
5 Limits on Scale Reliant Inference
@@
Having formally defined SRI as a problem of identifiability, we can prove fundamental limits on the types of inference that can be performed using data with arbitrary scale. We frame these limits in terms of common Frequentist and Bayesian inferential criteria. We show that there are strict limits on Frequentist inference but none on Bayesian inference. Still, we show that rigorous Bayesian inference requires a uncommon, specialized type of model called a Bayesian PIM. We use the concept of unacknowledged bias to discuss the practical implication of these limits.
5.1 Limits on Frequentist Inference
@@
The following four theorems follow naturally from the definition of SRI as a problem of identifiability. Similar impossibility proofs can be found in other fields that encounter unidentified models (e.g., Gabrielsen, (1978); Dufour, (1997); Gleser and Hwang, (1987)). Formal proofs of each theorem are provided in Supplementary Section B.3.
Theorem 1 (Consistency).
[…]
Theorem 2 (Unbiasedness).
[…]
Theorem 3 (Calibration).
@@
Suppose there exists $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). There does not exist a non-trivial [^1] Formally defined in Supplementary Section B.3, used to exclude trivial confidence sets that cover the entire parameter space. confidence set $S(\mathbf{Y})$ calibrated at level $01-\alpha>$ 0.
Theorem 4 (Type-I Error Control).
@@
Suppose there exists some $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). Then, for any $\alpha>0$, there does not exist a non-trivial [^2] Formally defined in Supplementary Section B.3, used to exclude trivial tests that never reject the null hypothesis. level-$\alpha$ test of the null hypothesis $\theta\in\Theta_{0}\subsetneq\Theta$.
To the best of our knowledge, these are the first formal proofs of what can and cannot be done when analyzing data with arbitrary scale. These theorems provide formal justification for CoDA’s insistence on scale invariance, yet they also show that that existing arguments within the CoDA community are overly restrictive. For example, consistent with CoDA’s arguments, these theorems show that the Frequentist ideal is not attainable for Full SRI problems, a non-trivial result. This contradicts a growing body of literature which claims to have developed rigorous methods for performing DA analysis from compositional data (e.g., Srinivasan et al., (2020); Love et al., (2014); Fernandes et al., (2014)). As we discussed in the prior section, this is a Full SRI problem. Theorem 2 proves that this is not possible. In Supplementary Section C, we review one such paper and discuss how to reconcile those claims with our results. Yet, these theorems also show that rigorous Frequentist inference is still possible in other types of SRI problems, namely Partial SRI problems. In fact, we have recently developed Frequentist methods for Differential Set Analysis (McGovern et al.,, 2023). In short, we have proven that there are some types of scale reliant estimation tasks that can be rigorously performed (relative to Frequentist ideals) using compositional data and demonstrated this claim in McGovern et al., (2023). In the next subsection, we expand beyond Frequentist inference and show that Bayesian inference is possible for all forms of SRI and SII.
5.2 Limits on Bayesian Inference
@@
The previous section established strict limits on Frequentist inference in the setting of SRI. There are no equivalent limits on Bayesian inference. So long as the chosen likelihood satisfies the likelihood principle and the prior is calibrated to prior beliefs, Bayesian inference is possible. Yet, for SRI, satisfying these two criteria can be challenging and requires a specialized type of model called a Bayesian PIM.
@@
For concreteness, consider LFC estimation from a full compositional survey $\mathbf{Y}$ (Full SRI). Suppose our prior beliefs can be represented as a non-singular distribution $p(\theta=(\theta_{1},\dots,\theta_{D}))$. Using Equation 2, this prior can alternatively be written as $p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$ where $\theta^{\tiny\mbox{comp}}$ is a vector with elements $\theta^{\tiny\mbox{comp}}_{d}$ which are functions of $\mathbf{W}^{{\tiny\mbox{comp}}}$ while $\theta^{\tiny\mbox{tot}}$ is a scalar that is a function of $W^{{\tiny\mbox{tot}}}$. To satisfy the likelihood principle, we must choose a likelihood model that accurately reflects the evidence of the observed data $\mathbf{Y}$ relevant to the model parameters $\theta$. Since $\mathbf{Y}$ provides no information about $W^{{\tiny\mbox{tot}}}$, our likelihood must satisfy
@@
$\mathcal{L}(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}};\mathbf{Y})\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};\mathbf{Y}).$
This likelihood, combined with this prior, is an example of a Bayesian PIM: it is a model in which the likelihood, even asymptotically, does not uniquely identify the parameters.
5.3 So What? Unacknowledged Bias
Considering that all models are wrong (Box,, 1976), it is natural to question the practical importance of the above limits: what is so bad about using a Bayesian model that is not a Bayesian PIM? To facilitate this discussion, we borrow the concept of unacknowledged bias from the Bayesian PIM literature (Gustafson,, 2015, pg. 35). Unacknowledged bias is a Frequentist concept which we, like prior literature (e.g., Gustafson, (2015)), choose to apply to both Bayesian and Frequentist methods. Like prior authors, we do not attempt to formally define unacknowledged bias, instead considering it only as a conceptual tool.
Suppose we had chosen a likelihood that was fully identified, e.g.,
$\mathcal{L}(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}};\mathbf{Y})\not\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};\mathbf{Y}).$
Assuming standard regularity conditions (e.g., $p(\theta)>0$ for all $\theta\in\Theta$), the posterior will concentrate to a point mass centered on the unique maximum likelihood estimate asymptotically. Although the model implies complete certainty in the estimate (zero variance), we have already established that the maximum likelihood estimate is incorrect (inconsistent and biased). This combination, an estimator that is biased yet fails to account for the bias through its variance, is called unacknowledged bias (Gustafson,, 2015). A hallmark of unacknowledged bias is a Type-I or Type-II error rate that increases with more data (see Supplemental Figure 2 for visual explanation). It is difficult to do rigorous science when using an estimator with unacknowledged bias. To use such an estimator, we must consider whether we have collected too much data to trust the results of our analysis. In Supplemental Section D, we discuss three common pitfalls encountered when developing Bayesian models for SRI including the use of Poisson or Negative Binomial likelihoods. We provide examples of these pitfalls within the existing literature and discuss how they lead to unacknowledged bias by violating either the likelihood or prior calibration principle.
In SRI, we cannot avoid bias (Theorem 2). Remarkably, Bayesian PIMs can acknowledge their bias and thereby minimize spurious scientific conclusions. Returning to our LFC example, consider that the posterior of the Bayesian PIM will factor as
$p(\theta^{\tiny\mbox{tot}},\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$ $=p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}},\mathbf{Y})p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$
$=p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y}).$
In short, the conditional relationship between $\theta^{\tiny\mbox{tot}}$ and $\theta^{\tiny\mbox{comp}}$ is the same before and after having seen the data. As a result, any learning that might occur regarding $\theta^{\tiny\mbox{tot}}$ after observing $\mathbf{Y}$ occurs indirectly through our updated beliefs about $\theta^{\tiny\mbox{comp}}$. So long as the variance of $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ is non-zero, the posterior will not concentrate to a point mass asymptotically. Instead, the posterior will acknowledge our lack of information about $\theta^{\tiny\mbox{tot}}$ and retain non-zero variance. This feature of Bayesian PIMs is why they are said to display an acknowledged rather than unacknowledged bias (Gustafson,, 2015). In Section 7, we demonstrate the practical advantages of acknowledging bias in the context of simulated and real data.

### v5

@@ 2.4 The Continuum between SII and SRI
Identifiability is not a dichotomous state but rather a continuum based on the identification region of a parameter. This implies that there is a spectrum between SRI and SII. Here, we define this spectrum before showing how these definitions apply to inferential problems described in existing literature.
@@
In the extreme case where ${\mbox{Id}}_{\theta}(P)$ is a singleton set, $\theta$ is identified, and the estimation problem is SII. At the opposite end where ${\mbox{Id}}_{\theta}(P)=\Theta$, $\theta$ is fully unidentified, and the estimation problem is Full SRI. Full SRI represents the extreme case where, due to the arbitrary scale of the data, all values of $\theta\in\Theta$ are consistent with the observed measurements. In between these two cases, where ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$, $\theta$ is partially identified, and the estimation problem is Partial SRI.
@@
In Section 2.5, we show that the distinction between SII, Full SRI, and Partial SRI is directly related to the conditions under which certain types of inference (e.g., Frequentist versus Bayesian) are possible. Before providing those results, we first motivate that all three types of estimation problems are encountered within the literature. To do this, we describe the two relevant factors that determine whether a problem is SII, Full SRI, or Partial SRI. The factors are the degree to which the target estimand depends on scale and the size of the identification region of $W^{{\tiny\mbox{tot}}}$ at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$.
2.4.1 Characterization of Target Estimands
Target estimands can be characterized based on the degree to which they depend on scale – whether they are scale invariant, partially scale reliant, or fully scale reliant. We define this characterization before providing examples of each type.
Let $\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$ denote the set of all systems that have a particular composition $\mathbf{W}^{{\tiny\mbox{comp}}}$. $\mathcal{W}^{\tiny\mbox{tot}}$ is the space of scale vectors, $\mathbb{Z}^{N+}$ if $W$ is integer valued or $\mathbb{R}^{N+}$ if continuous. Let $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$) denote the image of this set with respect to the target estimand:
$\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})=\{\theta\in\Theta:\exists\,\,W^{{\tiny\mbox{tot}}}\in\mathcal{W}^{\tiny\mbox{tot}},\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})=\theta\}.$
We define scale invariance/reliance based on the size of this image. If $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})$ is a singleton set, we say that $\theta$ is scale invariant at $\mathbf{W}^{{\tiny\mbox{comp}}}$; otherwise, it is scale reliant at $\mathbf{W}^{{\tiny\mbox{comp}}}$. We sub-categorize scale reliant estimands as fully scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))=\Theta$ or partially scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))\subsetneq\Theta$. Our definition reflects the following intuition. If $\theta$ is scale invariant, then its value is uniquely determined by $\mathbf{W}^{{\tiny\mbox{comp}}}$ alone and does not depend on $W^{{\tiny\mbox{tot}}}$. In contrast, if $\theta$ is partially scale reliant then $\mathbf{W}^{{\tiny\mbox{comp}}}$ limits the values that $\theta$ can take on but does not uniquely determine its value. If $\theta$ is fully scale reliant, then $\mathbf{W}^{{\tiny\mbox{comp}}}$ places no limits on the values $\theta$ can take on.
The scale reliance or scale invariance of a target estimand can depend on the composition $\mathbf{W}^{{\tiny\mbox{comp}}}$. For example, in a companion article, we have studied a set of non-parametric estimands (based on modified Kolmogorov-Smirnov statistics) which is scale invariant for some $\mathbf{W}^{{\tiny\mbox{comp}}}$ and scale reliant for others (McGovern, Nixon and Silverman, 2023). For simplicity, this manuscript only considers target estimands where their scale reliance or scale invariance holds for all $\mathbf{W}^{{\tiny\mbox{comp}}}$.
Example (Microbiome Data Analyses).
The LFC estimand described in Equation (2) is fully scale reliant. This can be established as follows. For any $\mathbf{W}^{{\tiny\mbox{comp}}}$ (and therefore any $\theta^{\tiny\mbox{comp}}_{d}\in(-\infty,\infty)$) and any $\theta_{d}\in(-\infty,\infty)$; there exists some $W^{{\tiny\mbox{tot}}}\in\mathcal{W}^{\tiny\mbox{tot}}$, and corresponding $\theta^{\tiny\mbox{tot}}\in(-\infty,\infty)$, such that $\theta_{d}=\theta^{\tiny\mbox{comp}}+\theta^{\tiny\mbox{tot}}$. That is, without restrictions on $\theta^{\tiny\mbox{tot}}\in(-\infty,\infty)$, fixing the value of $\theta^{\tiny\mbox{comp}}_{d}$ places no restrictions on the value of $\theta_{d}$.
At the opposite extreme, the types of log-ratio estimands purported by the field of compositional data analysis are scale invariant. For example, CoDA often analyzes log-ratio abundances between pairs of taxa $d_{i}$ and $d_{j}$ (Lovell et al., 2015):
$\eta_{n}=\log\frac{\mathbf{W}_{d_{i}n}}{\mathbf{W}_{d_{j}n}}.$ (4)
These quantities can be calculated from the system composition alone:
@@ 2.4.1 Characterization of Target Estimands
In between these two extremes there are partially scale reliant estimands. For example, in the analyse of human gut microbiome data a common task is to identify pairs of taxa whose abundances covary on the log-scale:
$\Sigma_{d_{i},d_{j}}=\text{cov}(\log\mathbf{W}_{d_{i}\cdot},\log\mathbf{W}_{d_{j}\cdot}).$ (6)
These covariance estimands underlie the construction of ecological interaction networks (Kurtz et al., 2015; Friedman and Alm, 2012). In Supplementary Section B.1, we prove this estimand is partially scale reliant: $\mathbf{W}^{{\tiny\mbox{comp}}}$ provides a finite lower bound for this estimate but it is unbounded from above regardless of $\mathbf{W}^{{\tiny\mbox{comp}}}$.
@@ 2.4.2 Characterization of Compositional Surveys
We categorize compositional surveys as partial or full compositional surveys depending on whether $W^{{\tiny\mbox{tot}}}$ is fully or partially identified by the model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$. That is, partial compositional surveys provide some information about the system scale ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)\subsetneq\mathcal{W}^{\tiny\mbox{tot}}$ whereas full compositional surveys provide no information ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)=\mathcal{W}^{\tiny\mbox{tot}}$. For brevity, this article focuses on full compositional surveys. Still, we highlight this distinction as we believe that further study of partial compositional surveys would be beneficial.
Example.
Authors who claim that microbiome data are compositional (e.g., Gloor et al. (2017)) are essentially stating that the data provides so little information about $W^{{\tiny\mbox{tot}}}$ that can be equated with simplicial vectors. We interpret such claims as a belief that the data comes from a full compositional survey. Empirical studies support this view illustrating that there is little relationship between the system scale and the scale of the data (Vandeputte et al., 2017; Props et al., 2017). While we expect that future researchers will identify specific cases where microbiome data does provide some degree of information about scale; for simplicity, this article models microbiome data as full compositional surveys.
Outside of microbiome studies there are clear cases where the data comes from partial compositional surveys. Consider political polling. Under standard conditions, a poll of 1000 voters in a town implies that the town must have at least 1000 voters. Still these data typically do not provide sufficient information to determine the true number of voters.
2.4.3 Theoretical Tools for Establishing SRI
In Supplemental Section B.2 we provide theoretical results which establish how estimands and compositional surveys combine to determine if an estimation problem is Full SRI, Partial SRI, or SII. We use those results to draw the following conclusions about estimation from full compositional surveys (e.g., microbiome data). LFC estimation (Eq 2, e.g., differential abundance analysis) is Full SRI. Covariance estimation (Eq. 6) is Partial SRI. Estimation of log-ratio quantities (e.g., Eq. 4) is SII.
2.5 Limits on Scale Reliant Inference
@@
Having formally defined SRI as a problem of identifiability, we can prove fundamental limits on the types of inference that can be performed using data with arbitrary scale. We frame these limits in terms of common Frequentist and Bayesian inferential criteria. We show that there are strict limits on Frequentist inference but none on Bayesian inference. However, rigorous Bayesian inference requires an uncommon type of model called a Bayesian Partially Identifiable Model (PIM). We use the concept of unacknowledged bias to discuss the practical implication of these limits.
2.5.1 Limits on Frequentist Inference
@@
We considers Frequentist inferential procedures motivated by canonical criteria such as consistency, unbiasedness, interval calibration, and type-I error control. The following four theorems follow from the definition of SRI as a problem of identifiability. Related impossibility proofs can be found in other fields that encounter partially identified models: such as econometrics (Dufour, 1997; Gabrielsen, 1978) or in the study of error-in-variable models (Gleser and Hwang, 1987). Proofs are provided in Supplementary Section B.3.
Theorem 2.2 (Consistency).
If ${\mbox{Id}}_{\theta}(P)$ is not a singleton set (Full or Partial SRI), then there does not exist a sequence of estimators for the target estimand $\theta$ that is consistent.
Theorem 2.3 (Unbiasedness).
If ${\mbox{Id}}_{\theta}(P)$ is not a singleton set (Full or Partial SRI), then there does not exist an estimator for the target estimand $\theta$ that is unbiased.
Theorem 2.4 (Calibration).
@@
Suppose there exists $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). There does not exist a non-trivial [^1] See Supplementary Section B.3. confidence set $S(\mathbf{Y})$ calibrated at level $01-\alpha>$ 0.
Theorem 2.5 (Type-I Error Control).
@@
Suppose there exists some $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). Then, for any $\alpha>0$, there does not exist a non-trivial [^2] See Supplementary Section B.3. level-$\alpha$ test of the null hypothesis $\theta\in\Theta_{0}\subsetneq\Theta$.
To the best of our knowledge, these are the first formal statements of what types of inferential criteria can or cannot be attained when analyzing data with arbitrary scale. These theorems provide formal justification for CoDA’s insistence on scale invariance, yet also show that existing arguments within the CoDA community are overly restrictive.
Consistent with CoDA’s arguments, these theorems show that the Frequentist ideal is not attainable for Full SRI problems. For example, Theorem 2, contradicts a growing body of literature which claims to have developed rigorous hypothesis tests for differential analysis of compositional data (Full SRI) that control Type-I error (Srinivasan, Xue and Zhan, 2020; Brill, Amir and Heller, 2022; Love, Huber and Anders, 2014; Fernandes et al., 2014). In Supplementary Section C, we review one paper that claims to have developed such a test and discuss how to reconcile those claims with our results.
Still, our results show that CoDA’s insistence on scale invariance goes to far. For example, our results show that non-trivial hypothesis tests could be developed for Partial SRI tasks. In fact, we have already done so in McGovern, Nixon and Silverman (2023). Moreover, if we look beyond Frequentist inference there are even more possibilities. In the next subsection, we show that Bayesian inference is possible for all forms of SRI.
2.6 Bayesian PIMs for SRI
@@
The previous section established strict limits on Frequentist inference in the setting of SRI. That is, limits on inferential procedures motivated by common criteria such as consistency. In contrast, this section explains why there are no equivalent limits on Bayesian inference as standard Bayesian assumptions can be satisfied. There are two principle assumptions which we review below: the likelihood principle and the prior calibration principle. While we explain how these assumptions can be satisfied, we also show that doing so requires a specialized type of model called a Bayesian PIM.
@@
Consider the problem of estimating LFCs for $D$ taxa ($\theta=(\theta_{1},\dots,\theta_{D})$) from a microbiome study $\mathbf{Y}$ (Full SRI). To satisfy the prior calibration principle, the prior must accurately represent prior beliefs about $\theta$. In most real-world scenarios we expect this would require a non-singular distribution $p(\theta)$, e.g., $\text{LogNormal}(\theta\mid\mu,\Sigma)$. Using Equation 3, such priors can reparameterized as $p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$ where $\theta^{\tiny\mbox{comp}}=(\theta^{\tiny\mbox{comp}}_{1},\dots,\theta^{\tiny\mbox{comp}}_{D})$ is a vector-valued function of $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta^{\tiny\mbox{tot}}$ is a scalar function of $W^{{\tiny\mbox{tot}}}$. To satisfy the likelihood principle, the chosen likelihood must accurately reflect the evidence provided by the observed data $\mathbf{Y}$ relevant to the model parameters $(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$. However, since the observed data comes from a full compositional survey, satisfying the likelihood principle requires a likelihood invariant to $W^{{\tiny\mbox{tot}}}$:
$\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}};\mathbf{Y})\propto\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}};\mathbf{Y}).$
As an example, when analyzing multivariate count data, a product multinomial might be appropriate:
@@
$\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}};\mathbf{Y})=\prod_{n}\text{Multinomial}(\mathbf{Y}_{\cdot n}\mid\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}).$
For LFC estimation, since $\theta^{\tiny\mbox{comp}}$ is solely a function of $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta^{\tiny\mbox{tot}}$ is solely a function of $W^{{\tiny\mbox{tot}}}$, this could be reparameterized as a likelihood of the form $\mathcal{L}(\theta^{\tiny\mbox{comp}},\lambda;\mathbf{Y})$ where $\lambda$ is some nuisance parameter that depends on $\mathbf{W}^{{\tiny\mbox{comp}}}$. For brevity, we omit $\lambda$ from future notation. This combination of non-singular prior $p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$ with a likelihood that is invariant to one of the prior parameters $\mathcal{L}(\theta^{\tiny\mbox{comp}};\mathbf{Y})$ is an example of a Bayesian PIM. A more general introduction to Bayesian PIMs can be found in Gustafson (2015).
Bayesian PIMs can have posteriors that demonstrate appealing, yet non-traditional behavior. Many traditional Bayesian models have posteriors which concentrate to a point-mass asymptotically (Schwartz, 1965). In SRI, this behavior is undesirable. From Theorem 2.3, any point estimator is biased and therefore any Bayesian posterior that asymptotically collapses to a point mass is problematic. Such methods are said to display an unacknowledged bias as they are biased yet fail to account for that bias in their estimation of uncertainty (Gustafson, 2015, pg. 35). Unlike those traditional methods, Bayesian PIMs often have non-zero asymptotic variance. More concretely, consider that the posterior for the above model factors as:
$p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}}\mid Y)$ $\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};Y)p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$
$\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};Y)p(\theta^{\tiny\mbox{comp}})p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$
$\propto p(\theta^{\tiny\mbox{comp}}\mid Y)p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}}).$ (7)
So long as the prior is non-singular, the posterior factor $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ will not collapse, even asymptotically. The value of a posterior that maintains non-zero variance is straight-forward: if the posterior is biased, then only a posterior that has non-zero variance has any chance of covering the true value. It is for this reason that Bayesian PIMs are often said to display an acknowledged rather than unacknowledged bias. Reinforcing this terminology, one author has even shown that the posterior of Bayesian PIM’s can be interpreted as a probabilistic model for the bias of the posterior mean (Gustafson, 2015, pg. 35).
A major motivation of this work is to avoid unacknowledged bias. To underscore the danger, consider that a hallmark of unacknowledged bias is a Type-I or Type-II error rate that increases with more data (see Supplemental Figure 2 for visual explanation). This behavior is counter to standard statistical wisdom – where statistical procedures are supposed to improve with more data. For concreteness, in Supplemental Section D we discuss three common pitfalls encountered when developing Bayesian models for SRI including the use of Poisson or Negative Binomial likelihoods. That section discusses how each pitfall can lead to unacknowledged bias. In Section 4, we demonstrate the practical benefits of Bayesian PIMs for microbiome data analysis using both simulated and real data studies. Those studies demonstrate that common methods for microbiome differential abundance analysis demonstrate unacknowledged bias. They also show how Bayesian PIMs can drastically reduce Type-I and even Type-II error rates by acknowledging their own bias.